11 3 -21 10. Use properties of determinants to find the value of the determinant 6 1 -5 18 2 3
Answer #1

IC co -51 6. 6 1 -5 3 + (-2) 2 8 3 0 2 3 2 3 1 [3x1 = (-5)x2] 3f6x3-4-54x8] -26x2=8*t] - (3 + 10) -3 (18+ 40) - 2 (12-8) - 2x